OpenStudy (anonymous):

Need assistance in "Solving and Graphing Systems of Equations"... Find the solution of this system of equations. 6y = -50-4x 5x + 6y = -49

OpenStudy (wolf1728):

A) 4x + 6y = -50 B) 5x + 6y = -49 Multiply B by -1 B) -5x -6y = 49 then add this to A A) 4x + 6y = -50 -x = -1 x = 1 inputting x = 1 into equation A 4 + 6y = -50 6y = -54 y = -9

OpenStudy (anonymous):

Okay, thanks!!!! So the finalized answer is -1,-9?

OpenStudy (wolf1728):

No it is x =1 y= -9

OpenStudy (anonymous):

Oh oops. Sorry so 1,-9?

OpenStudy (wolf1728):

Yes

OpenStudy (anonymous):

Okay, would I use brackets like this? {1,-9}

OpenStudy (wolf1728):

Or I think parentheses are also okay (1, -9)

OpenStudy (anonymous):

Okay, do you mind if I tag you in others? :)

OpenStudy (wolf1728):

In other questions? okay but I'm getting away from the computer at the moment

OpenStudy (anonymous):

Alright, how long until you're back? :)

OpenStudy (wolf1728):

Could be hours.

OpenStudy (anonymous):

Dang.. Ok.

OpenStudy (wolf1728):

maybe I could answer 1 more question :-)

OpenStudy (wolf1728):

you do have another question right?

OpenStudy (anonymous):

Should I post here?

OpenStudy (wolf1728):

sure - that would be fine

OpenStudy (anonymous):

Find the solution of this system of equations. -7x + 6y = 59 -8x + 6y = 58

OpenStudy (wolf1728):

A) -7x + 6y = 59 B) -8x + 6y = 58 multiply A by -1 A) 7x -6y = -59 then add to B B) -8x + 6y = 58 -x = -1 x =1 -8x + 6y = 58 -8 + 6y = 58 6y = 66 y = 11

OpenStudy (anonymous):

Thanks! :D

OpenStudy (wolf1728):

u r welcome :-)